OpenStudy (anonymous):

[7.04] 1.Create the division of two monomials by filling in the blanks with exponents greater than one. a to the power of ? over a to the power of ? 2.Find the quotient of these monomials. 3.Explain which operation was done on the exponents to find this quotient. 4.Prove this operation provides the correct quotient by showing the fraction in expanded form.

OpenStudy (anonymous):

Please, I need help, and lots of it. :'(

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

Okay...so say we have \[\huge \frac{a^6}{a^2}\]

OpenStudy (johnweldon1993):

When you are dividing exponents with the same base number (a here) you just subtract the exponents so \[\huge \frac{a^6}{a^2} = a^{6 - 2} = a^4\]

OpenStudy (johnweldon1993):

And to show it...we have \[\huge \frac{\cancel{a} \times \cancel{a} \times a \times a \times a \times a}{\cancel{a} \times \cancel{a}}\] As you cansee...there are only 4 a's left...so our answer is indeed \(\huge a^4\)

OpenStudy (johnweldon1993):

Make sense? :)

OpenStudy (anonymous):

3.Explain which operation was done on the exponents to find this quotient.? What was the answer for this?

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

Subtraction was done...that is where we subtracted 2 from 6 :) @mikaela.mellish